There are 3 assets. The following information is available about the assets:

Asset

Expected return

Expected standard deviation

Beta

A

.10

.04

1.1

B

.14

.06

1.5

C

.04

0.0

0

a. (5 points) What is the expected return on a portfolio that is 1/3 invested in each asset?

b. (5 points) What is the Beta of a portfolio that is 50% A and 50% C?

c. (5 points) If you wanted to have a portfolio with a Beta of 1.04 and had 20% in asset C, what percentages would you invest in A and B?
Answer #1

Answer A)

Expected Return = W(A) * R(A) + W(B) * R(B) + W(C) * R(C)

= 1/3* 10% + 1/3 * 14% + 1/3 * 4%

= 9.33%  

Answer B) Beta is weighted Average

Beta of Portfolio with A & C = 50% * 1.1 + 50% * 0

= 0.55

Answer C)

let weight of A is X, then Weight of B becomes 1 - X - 0.2

beta of portfolio =  1.1 * X + 1.5* (1-X-0.2) + 0 * 20%

1.04 = 1.1X + 1.5 - 1.5 X - 0.3

1.04 = 1.2 - 0.4 X

X = (1.2 - 1.04) / 0.4

X = 40%

Invest in A = 40%

Invest in B = 100 - 40%-20% = 40%

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.